OpenStudy (anonymous):

solve each system by substitution y=-5x+22 -7x+5y=14

OpenStudy (anonymous):

Slow down there. Other people need help too. Substitute the entire first equation into the second one where there is a y. You then have an equation in one variable. Solve for x. Go back to the first equation with your x-value and solve for y

OpenStudy (anonymous):

you just need to plug y=-5x+22 to the equation -7x+5y=14 -7x+5(-5x+22)=14 solve that and you'll get the value of x and y

OpenStudy (anonymous):

y=-5x+22 -----------1 -7x+5y=14------------2 put 1 into 2, -7x+5y=14 -7x+5(-5x+22)=14 -7x+25x+110=14 -18x+110=14 -18x=-96 x=-96/-18---------3 put 3 into 1, y=-5x+22 y=5(-96/-18)+22 y=48/2/3